### v1

# DART-LLM: Dependency-Aware Multi-Robot Task Decomposition and Execution using Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated significant reasoning capabilities in robotic systems. However, their deployment in multi-robot systems remains fragmented and struggles to handle complex task dependencies and parallel execution. This study introduces the DART-LLM (Dependency-Aware Multi-Robot Task Decomposition and Execution using Large Language Models) system, designed to address these challenges. DART-LLM utilizes LLMs to parse natural language instructions, decomposing them into multiple subtasks with dependencies to establish complex task sequences, thereby enhancing efficient coordination and parallel execution in multi-robot systems. The system includes the QA LLM module, Breakdown Function modules, Actuation module, and a Vision-Language Model (VLM)-based object detection module, enabling task decomposition and execution from natural language instructions to robotic actions. Experimental results demonstrate that DART-LLM excels in handling long-horizon tasks and collaborative tasks with complex dependencies. Even when using smaller models like Llama 3.1 8B, the system achieves good performance, highlighting DART-LLM’s robustness in terms of model size. Please refer to the project website https://wyd0817.github.io/project-dart-llm/ for videos and code.
I Introduction
Large Language Models (LLMs) have demonstrated significant reasoning capabilities that can be applied to robot systems. Chen et al. introduced the Decision Transformer, a reinforcement learning method that redefines decision-making processes as sequence modeling problems [1]. Reed et al. developed Gato, a generalist artificial intelligence model capable of learning and adapting across multiple tasks and modalities [2]. Ahn et al.’s FLAN-SayCan and PaLM-SayCan leverage the capabilities of large pre-trained language models (LLMs) to interpret and execute tasks based on natural language instructions, enabling robots to understand and perform tasks driven by complex natural language commands [3]. Huang et al.’s Inner Monologue method integrates language models into planning and reasoning processes, enhancing robots’ decision-making capabilities [4]. Brohan et al. proposed RT-1, a model for real-time control and decision-making in complex environments, applying transformer architecture directly to robotics technology to ensure unprecedented precision and adaptability in understanding and navigating the physical world [5]. Brohan et al. further introduced the RT-2 model, which integrates vision, language, and action capabilities, enabling robots to autonomously learn and control actual actions from web knowledge [6]. Driess et al.’s Palm-E model enhances the perception and action capabilities of language models, offering a new way for artificial intelligence systems to interact more naturally with the external world through enhanced multimodal understanding [7]. Liang et al. launched Code as Policies, enabling large language models (LLMs) to convert natural language commands into robot strategy code with minimal prompts [8]. Huang et al. introduced VoxPoser, which extracts language-conditioned affordances and constraints from LLMs and applies them to the perceptual space through VLMs. VoxPoser uses a code interface without additional training for any component [9].
Multi-robot systems have shown immense potential across various scenarios, such as complex construction sites and disaster response [10], due to their high flexibility and efficiency. The key advantage of these systems lies in their ability to perform tasks in parallel and collaborate to solve problems that are difficult for a single robot to handle independently. Zhao et al. proposed the RoCo framework, utilizing LLMs for high-level communication and low-level path planning [11]. Kannan et al. developed the SMART-LLM framework, specifically targeting task planning issues in multi-robot systems [12]. The comparison of the above-mentioned studies is shown in Table I. Table I. indicates that although existing research has significantly enhanced robots’ understanding of natural language and flexibility in task execution, driven by cutting-edge technologies like LLMs, there remain fragmented deployments of multi-robot systems. Moreover, these studies still face limitations in handling complex task dependencies and parallel execution. They utilize linear methods for task decomposition, making it challenging to effectively manage the interdependencies among subtasks generated from tasks with complex dependencies, which is particularly critical in multi-robot systems. The key advantage of multi-robot systems lies in their ability to collaboratively solve problems that are difficult for a single robot to handle independently. However, achieving this collaboration requires an effective mechanism that ensures logical sequencing of interdependent subtasks among multiple robots, while enabling parallel execution of independent tasks. To address these issues, we propose the DART-LLM (Dependency-Aware Multi-Robot Task Decomposition and Execution using Large Language Models) system. The main contributions are as follows:
1. Integrated System Framework: An integrated system framework named DART-LLM, comprising multiple modules, is proposed. The framework includes the QA LLM module, Breakdown Function module, Actuation module, and a Vision-Language Model (VLM)-based object detection module. These modules perform Instruction Parsing and Task Decomposition, Parsing of Decomposed Tasks, Grounding in Embodiments, and Updating the Object Map Database to accomplish task decomposition and execution from natural language instructions to robotic actions.
2. Dependency-Aware Task Decomposition: A dependency-aware task decomposition method is proposed, which breaks down complex tasks into multiple interdependent subtasks. By accounting for dependencies, the DART-LLM allows execute dependent subtasks sequentially while running independent tasks in parallel, thereby enhancing logical capabilities and improving the efficiency of complex tasks.
TABLE I: Comparison of Related Work in Multi-Robot Systems and Large Language Models (LLMs)
Related Work Multi-Robot System Robot Mobility Real-time Capability Dependency-Aware
Chen et al., Decision Transformer [1]
Reed et al., Gato [2]
Ahn et al., FLAN-SayCan, PaLM-SayCan [3]
Huang et al., Inner Monologue [4]
Brohan et al., RT-1 [5]
Brohan et al., RT-2 [6]
Driess et al., Palm-E [7]
Liang et al., Code as Policies [8]
Huang et al., VoxPoser [9]
Zhao et al., RoCo [11]
Kannan et al., SMART-LLM [12]
Proposed DART-LLM
Figure 1: An overview of the DART-LLM system architecture. The system is divided into three main modules: the Sensor Module, the Intelligent Command Interface Module, and the Actuation Module. The Sensor Module captures and processes data using a Vision-Language Model (VLM)-based object detector, updating the object map with detected items. The Intelligent Command Interface Module interprets user instructions via a Question-Answering Large Language Model (QA LLM), decomposing tasks into subtasks with dependencies through the Breakdown Function Parser and Handler. This allows the establishment of complex task dependencies and coordination between multiple robots. Finally, the Actuation Module executes real-time operations using the ROS Navigation stack, guiding each robot according to the parsed and dependency-aware instructions.
II Problem Formulation
This section presents the foundational mathematical concepts of the DART-LLM system, including the definitions of robot sets, associated skills, task decomposition, team formation, and navigation functions essential for task execution.
II-A Robot Set and Skills
We define the set of all robots as:
=$\mathbb{R}=\{R^{1},R^{2},\ldots,R^{N}\}$ (1)
where $N$ is the number of robots. The set of skills for each robot is:
$\mathbb{S}=\{S^{1},S^{2},\ldots,S^{N}\}$ (2)
where $S^{n}\subseteq\Delta$ and $\Delta$ is the set of all skills. Each robot $R^{n}$ has a skill set $S^{n}$.
II-B Team of Robots and Navigation Functions
A team of robots is defined as:
$\mathbb{A}=\{A^{1},A^{2},\ldots,A^{Q}\}$ (3)
where $Q$ is the number of robots in the team. The skills of a team of robots are:
$\mathbb{S}_{A}=\{S_{A}^{1},S_{A}^{2},\ldots,S_{A}^{Q}\}$ (4)
The $m$ navigaiton function skills of a robot are defined as:
$\mathbb{S}_{F}=\{S_{F}^{1},S_{F}^{2},\ldots,S_{F}^{m}\}$ (5)
The set of all navigation functions for all robots is defined as
$T_{S}^{f}\subseteq\bigcup\mathbb{S}_{F}$ (6)
The skills needed by the team must satisfy:
$T_{S}^{k}\subseteq\bigcup\mathbb{S}_{A}$ (7)
Thus, the combined skills must cover all sub-task and function requirements:
$\bigcup\mathbb{S}_{AF}\supseteq T_{S}^{k}\cup T_{S}^{f}$ (8)
II-C Task and Function Requirements
The high-level language instruction is represented as $\mathbb{I}$ and the environment is denoted by $\mathbb{E}$. The instruction $\mathbb{I}$ is decomposed using the QA LLM module into a set of sub-tasks. The details of this module will be described in subsection III-A:
$\mathbb{T}=\{T^{1}_{t_{1}},T^{1}_{t_{2}},\ldots,T^{K}_{t_{j}}\}$ (9)
where $K$ is the number of sub-tasks, and $t_{j}$ denotes the temporal order of a sub-task, with $j\leq K$. Each task $T^{K}_{t_{j}}$ should be executable using the skills from the skill library within $\bigcup\mathbb{S}_{AF}$.
II-D Dependency Relationships Between Tasks
Complex tasks in multi-robot systems often involve multiple subtasks with interdependencies. These dependencies are managed through task decomposition using a Directed Acyclic Graph (DAG). The DAG is defined as $G=(\mathbb{T},\mathbb{D})$, where:
$\mathbb{D}\subseteq\mathbb{T}\times\mathbb{T}$ (10)
represents the set of directed edges indicating dependencies between subtasks. An edge $(T^{k-1}_{t_{i}},T^{k}_{t_{j}})\in\mathbb{D}$ indicates that subtask $T^{k}_{t_{j}}$ depends on the successful completion of subtask $T^{k-1}_{t_{i}}$.
III The DART-LLM Framework
This section provides an overview of DART-LLM. The overall control architecture of DART-LLM is illustrated in Fig. 1. Communication between all modules is based on Robot Operating System 2 (ROS2) topics [13].
III-A Instruction Parsing and Task Decomposition Use the QA LLM Module
High-level natural language instructions $\mathbb{I}$ serve as the input for DART-LLM. The input is first processed by the QA LLM for Instruction Parsing and Task Decomposition. The QA LLM parses the high-level instructions $\mathbb{I}$ while considering the environment $\mathbb{E}$ and the skills $\bigcup\mathbb{S}_{AF}$ possessed by all robots $\mathbb{R}$. The skills in $\bigcup\mathbb{S}_{AF}$ are derived from all robots $\mathbb{R}$ and any possible teams $\mathbb{A}$ they may form. It should be noted that during Task Decomposition, DART-LLM uses awareness of dependency relationships $\mathbb{D}$ to decompose the task into multiple subtasks $\mathbb{T}=\{T_{t_{1}}^{1},T_{t_{2}}^{1},\ldots,T_{t_{j}}^{K}\}$. This facilitates parallel execution of decomposed subtasks without mutual dependencies and clearly expresses dependencies between subtasks. Therefore, the prompt $\mathbb{P}$ for the QA LLM is defined as $\mathbb{P}=(\mathbb{I},\mathbb{E},\mathbb{R},\mathbb{S}_{A},\mathbb{S}_{F},\mathbb{F})$ , where $\mathbb{F}$ represents the few-shot example. The output of the QA LLM uses a structured standard format to handle the parsed commands and their dependencies. The standardized structure follows JSON syntax, as shown below:
"instruction_function": {
"name": "<breakdown function 1>",
"dependencies":["<dep 1>", "<dep 2>", …, "<dep n>"]
"object_keywords": ["<key 1>", "<key 2>", …, "<key n>"],
"instruction_function": {
"name": "<breakdown function 2>",
"dependencies":["<dep 1>", "<dep 2>", …, "<dep n>"]
"object_keywords": ["<key 1>", "<key 2>", …, "<key n>"],
"instruction_function": {
"name": "<breakdown function m>",
"dependencies":["<dep 1>", "<dep 2>", …, "<dep n>"]
"object_keywords": ["<key 1>", "<key 2>", …, "<key n>"]
The structured standard format message contains the names of the breakdown functions (i.e., atomic action functions) along with their parameters, object keywords, and dependencies. These atomic action functions belong to $\mathbb{S}_{F}$ and $\mathbb{S}$. For specific atomic action functions, please refer to subsection III-C. The parameters of the atomic action functions are searched within the object map database. For details, please refer to subsection III-D.
III-B Parsing of Decomposed Tasks Uses the Breakdown Function Modules
The Breakdown Function Parser module receives the subtasks $\mathbb{T}=\{T_{t_{1}}^{1},T_{t_{2}}^{1},\ldots,T_{t_{j}}^{K}\}$ decomposed by the QA LLM Module and parses the instruction functions, parameters, and dependencies of all the skills used by the subtasks. The parsed content is then sent to the Breakdown Function Handler module. The Breakdown Function Handler module searches for the location information of entities using object keywords in the object map database. This database is updated using the VLM-based object detector module.
III-C Grounding in Embodiments Uses the Actuation Module
The Actuation module grounds in embodiments by mapping $\mathbb{T}=\{T_{t_{1}}^{1},T_{t_{2}}^{1},\ldots,T_{t_{j}}^{K}\}$ to $\bigcup\mathbb{S}_{AF}$. This is specifically achieved through predefined atomic action functions. All atomic action functions are asynchronously executed by the Actuation module. These atomic action functions are divided into two categories: navigation functions and robot-specific functions, corresponding to $\mathbb{S}_{F}$ and $\mathbb{S}$, respectively. The implementation of $\mathbb{S}_{F}$ is achieved through the ROS Navigation stack [14], while $\mathbb{S}$ is realized by creating specific action sets for each robot.
III-C1 Navigation Functions
The navigation functions are responsible for directing the movement and area access permissions of all robots or specific robots within a designated area. These functions include whether to avoid certain areas, return to the initial position, or move to a specified location. Table  II provides detailed descriptions of each navigation function.
TABLE II: Navigation Function Descriptions
Number Function Name Description
N1 avoid_areas_for_all_robots Sets the cost map to make all robots avoid specified areas.
avoid_areas_for_specific_robots Sets the cost map for selected robots to avoid specified areas.
N2 target_area_for_all_robots Guides all robots to target points near the specified area.
target_area_for_specific_robots Guides selected robots to target points near the specified area.
N3 allow_areas_for_all_robots Configures the cost map to allow all robots to access specified areas.
allow_areas_for_specific_robots Configures the cost map for selected robots to access specified areas.
N4 return_to_start_for_all_robots Instructs all robots to return to their initial starting position.
return_to_start_for_specific_robots Instructs selected robots to return to their initial starting position.
III-C2 Robot Functions
The robot-specific functions are tailored to the unique operational capabilities of different robot types, such as excavators and dump trucks. These functions handle tasks such as digging, unloading, and loading. Table  III provides detailed descriptions of each robot function. It should be noted that different types of robots have different atomic action function definitions.
TABLE III: Robot Function Descriptions
Number Function Name Description
FE1 excavator_digging Instructs the excavator to dig at the specified target location.
FE2 excavator_unloading Instructs the excavator to unload at the specified target location.
FD1 dump_loading Instructs the dump truck to load materials at the specified target location.
FD2 dump_unloading Instructs the dump truck to unload materials at the specified target location.
III-D Updating Object Map Database Uses the VLM-based object detector module
The DART-LLM updates the object map database in environment $\mathbb{E}$ using the VLM-based object detector module. This module first converts all images captured by Unmanned Aerial Vehicles (UAVs) into bird’s-eye views. These views are processed to detect and recognize objects in the environment $\mathbb{E}$ Selective search [15] is then used on the bird’s-eye views to generate candidate bounding boxes. Following region proposals, a CLIP-based model [16] is employed for identification, updating the object map database with the names, locations, and shapes of the objects.
IV Experiments
IV-A Experimental Setup
To evaluate the performance of the DART-LLM system across different task sets, this study uses construction robots, which are highly sensitive to the order of sub-tasks, as test subjects to assess the performance of DART-LLM. Errors in the sequence of sub-tasks may lead to experimental failure; for instance, if a dump truck departs the soil pile before the excavator completes unloading, the loading and unloading operation will fail. This study designed a test set comprising three task levels (L1, L2, L3) and six test scenarios. Each task level includes different tasks to evaluate the performance of various models in a simulated environment. The task numbering reflects its respective level.
The experimental environment is based on the Unity simulation platform. The simulator uses the PhysX physics engine. The simulation environment includes two C30R crawler transport robots produced by Japan’s Yanmar and one ZX120 excavator produced by Japan’s Hitachi Construction Machinery.
Table IV lists the specific tasks, required skills, and applicable robot types for each task level. Level 1 tasks mainly assess the ability of a single robot to perform simple tasks; Level 2 tasks require multiple robots to work cooperatively; Level 3 tasks are more complex, requiring robot teams to consider dependencies and the sequence of subtasks.
TABLE IV: Description of specific tasks in each task level
Task Level Task Number Task Description Skill Number Applicable Robots
Level 1 L1-T1 Inspect a puddle N1, N2, N3, N4 Excavator or Dump Truck
L1-T2 Clear an obstacle N1, N2, N3, N4, FE1, FE2 Excavator
Level 2 L2-T1 Excavate soil N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
L2-T2 Transport soil to the dump truck’s initial position N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
Level 3 L3-T1 Clear the obstacle, then dig soil N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
L3-T2 Clear the obstacle, then inspect the puddle N1, N2, N3, N4, FE1, FE2, FE2, FD2 Excavator, Dump Truck
IV-B Evaluation Metrics
To comprehensively assess the performance of different models in the tasks mentioned above, this study uses the following key metrics:
  • SR (Success Rate): Whether the task is completely successful.
  • IPA (Instruction Parsing Accuracy): The accuracy of instruction parsing, defined as the proportion of correctly parsed instructions mapped to the correct atomic action function names and parameters.
  • DSR (Dependency Satisfaction Rate): The rate of dependency satisfaction, referring to the proportion of subtasks completed in the correct order of dependencies.
  • SGSR (Semantic Grounding Success Rate): The rate of successful semantic grounding, defined as the proportion of code generated that can be successfully parsed by the breakdown function parse module.
IV-C Experimental Results
IV-C1 Testing on the Test Set Using Different LLMs
This study tested DART-LLM using GPT-4o, GPT-3.5-turbo, Llama3.1, and Claude3.5. Table V shows the evaluation results of different models on tasks of L1, L2, and L3. The results indicate that for task level L1, all models achieved perfect scores across all metrics (SR=1.00, IPA=1.00, DSR=1.00, SGSR=1.00). As task complexity increases (from L1 to L3), performance differences among models become apparent. In task level L2, the IPA scores for all models were 1. The SR of the Llama 3.1 model was 0.87, higher than GPT-3.5-turbo’s 0.83, influenced by SGSR. This may be attributed to Llama 3.1’s better instruction-following ability, resulting in the generation of compliant JSON commands. For GPT-3.5-turbo, the SR dropped to 0.83 in L2 tasks, with a DSR of 0.87 and SGSR of 0.83. In contrast, the GPT-4o model maintained a high level of performance with an SR of 0.97, and both DSR and SGSR scores of 0.97. This can be attributed to GPT-4o’s enhanced reasoning and instruction-following abilities, allowing it to think more logically and generate compliant commands, leading to higher DSR and SGSR scores. The Claude 3.5 model performed worse than GPT-4o but outperformed both the Llama 3.1 and GPT-3.5-turbo models. All models performed well in terms of IPA, which can be attributed to the introduction of the structured JSON format and dependency-aware task decomposition method, allowing models to better focus on the task decomposition itself. At task level L3 (the highest complexity), the GPT-4o model continued to excel, maintaining scores similar to those in task level L2, with an SR of 0.97, IPA of 1.00, and both DSR and SGSR scores of 0.97. Other models showed a decline in performance in L3 tasks, particularly GPT-3.5-turbo, whose SR dropped to 0.75 due to its insufficient instruction-following capability. Overall, the DART-LLM system showed optimal performance when using the GPT-4o model, especially in handling more complex tasks. Notably, the Llama 3.1 model, with only 8B parameters, performed better than GPT-3.5-turbo, despite not achieving the best results. In task level L3, while Llama 3.1’s DSR score was lower than that of GPT-3.5-turbo, its higher SGSR resulted in an SR that was superior to GPT-3.5-turbo. This indicates that using smaller models with higher instruction-following capabilities is the best choice for the DART-LLM system, as smaller models offer better deployability and real-time performance.
TABLE V: Evaluation results of different models on L1, L2, and L3 tasks
Models L1 L2 L3
SR IPA DSR SGSR SR IPA DSR SGSR SR IPA DSR SGSR
DART-LLM (Llama3.1)
1.00 1.00 1.00 1.00 0.87 1.00 0.87 0.93 0.83 0.93 0.83 0.89
DART-LLM (GPT-3.5-turbo)
1.00 1.00 1.00 1.00 0.83 1.00 0.87 0.83 0.75 0.97 0.87 0.75
DART-LLM (GPT-4o)
1.00 1.00 1.00 1.00 0.97 1.00 0.97 0.97 0.97 1.00 0.97 0.97
DART-LLM (Claude3.5)
1.00 1.00 1.00 1.00 0.89 1.00 0.89 0.97 0.89 1.00 0.89 0.93
IV-C2 Ablation Study
To investigate the impact of dependencies on the efficiency of multi-robot systems, we utilized the Llama3.1 model to test six task sets under conditions with and without the dependency-aware approach. The experimental results, as shown in Fig. 2, indicate that the dependency-aware approach significantly saves time in complex tasks. This is because the dependency-aware approach allows subtasks to be executed sequentially while also enabling parallel execution of subtasks without dependency relationships, indicating that the use of the dependency-aware approach enhances the system’s logical capabilities.
Figure 2: Comparison of Dependency-Aware and Non-Dependency-Aware Approaches across Various Tasks.
IV-C3 Deployment in Real World
To evaluate the deployability of DART-LLM, we tested it in a real-world setting. Fig. 3 shows the results of the L2-T1 task. The results indicate that DART-LLM demonstrates good applicability on actual robots. For more experiments, please visit the project website project-dart-llm.
Figure 3: DART-LLM (Llama3.1) in the L2-T1 Task Using Real Robots: The sequence begins at 0 [s] with the initial positioning of the Dump Truck at the starting location. At 8 [s], the Excavator prepares to dig soil. By 24 [s], the Excavator transfers the excavated soil into the Dump Truck, completing the loading operation. At 48 [s], the loaded Dump Truck navigates towards the designated dumping area. At 56 [s], the Dump Truck unloads the soil at the puddle location. Finally, at 64 [s], the Dump Truck completes the task by retracting its vessel.
V Conclusion and Future Work
This paper presents the DART-LLM, which leverages LLMs to achieve dependency-aware task decomposition and execution in multi-robot systems. By decomposing tasks into multiple subtasks with dependencies, DART-LLM effectively manages complex task sequences, facilitating parallel execution and collaborative cooperation in multi-robot systems. Experimental results validate the effectiveness of DART-LLM in handling complex tasks and dependencies, with the system showing optimal performance when using the GPT-4o model. Importantly, even when using smaller models like Llama 3.1 8B, DART-LLM demonstrated high success rates and efficiency, proving the system’s adaptability and robustness across different model sizes. Future work includes enhancing DART-LLM’s applicability to larger-scale robot teams.
ACKNOWLEDGMENT
This work is supported by JST [Moonshot Research and Development], Grant Number [JPMJMS2032]. This work was submitted to the 2025 IEEE International Conference on Robotics and Automation on September 15, 2024.